if every 9th person in line for concert tickets gets a free t shirt, and every 24th person gets a free ticket, which person will be the first to a t-shirt and free ticket? how do you know?
Found 2 solutions by checkley71, ankor@dixie-net.com:
Answer by checkley71(8403)

PERSON NUMBER 72 WILL GET BOTHA FREE TICKET & A T SHIRT.
9+9=18+9=27+9=36+9=45+9=54+9=63+9=72
24+24=48+24=72 THE FIRST MATCH.

Answer by ankor@dixie-net.com(22740)

f every 9th person in line for concert tickets gets a free t shirt, and every 24th person gets a free ticket, which person will be the first to a t-shirt and free ticket? how do you know?
:
Wouldn't it be the first number that is a multiple 9 and 24?
:
9 * 24 = 216 would be one, but is it the first one?
:
No, 72 would be the first one, 72 divided by 24 = 3. A prime number, it's the lowest.
: